# Global Behavior of Learning Dynamics in Zero-Sum Games with Memory Asymmetry

### v1

@@ Abstract
This study examines the global behavior of dynamics in learning in games between two players, X and Y. We consider the simplest situation for memory asymmetry between two players: X memorizes the other Y’s previous action and uses reactive strategies, while Y has no memory. Although this memory complicates the learning dynamics, we discover two novel quantities that characterize the global behavior of such complex dynamics. One is an extended Kullback-Leibler divergence from the Nash equilibrium, a well-known conserved quantity from previous studies. The other is a family of Lyapunov functions of X’s reactive strategy. These two quantities capture the global behavior in which X’s strategy becomes more exploitative, and the exploited Y’s strategy converges to the Nash equilibrium. Indeed, we theoretically prove that Y’s strategy globally converges to the Nash equilibrium in the simplest game equipped with an equilibrium in the interior of strategy spaces. Furthermore, our experiments also suggest that this global convergence is universal for more advanced zero-sum games than the simplest game. This study provides a novel characterization of the global behavior of learning in games through a couple of indicators.
1 Introduction
@@
Learning in games targets how multiple agents learn their optimal strategies in the repetition of games [1]. The set of such players’ best strategies is defined as Nash equilibrium [2], where every player has no motivation to change his/her strategy. However, this equilibrium is hard to compute in general because one’s best strategy depends on the others’ strategies. Indeed, the behavior of multi-agent learning is complicated in zero-sum games, where players conflict in their payoffs. Even when players try to learn their optimal strategies there, their strategies often cycle around the Nash equilibrium and fail to converge to the equilibrium.
@@
In order to understand such strange behaviors, which are unique in multi-agent learning, the dynamics of how multiple agents learn their strategies, say, learning dynamics, are frequently studied [3, 4, 5, 6]. The representative dynamics of interest are the replicator dynamics, which is based on the evolutionary dynamics in biology [7, 8, 9, 10, 11]. These dynamics are also known as the multiplicative weight updates (MWU) in its discrete-time version [12, 13]. Furthermore, their connection to other representative learning dynamics, such as gradient ascent [14, 15, 16, 17] and Q-learning [18, 19, 20], should be noted. Such replicator dynamics are known to be characterized by Kullbuck-Leibler (KL) divergence, which is the distance from the Nash equilibrium to the players’ present strategies. This KL divergence is conserved during the learning dynamics, and the distance from the Nash equilibrium is invariant. Follow the Regularized Leader (FTRL) is a class of learning algorithms including the replicator dynamics and also has its conserved quantity, which is the summation of divergences for all the players [21, 22]. To summarize, such complex learning dynamics have been discussed based on their conserved quantity.
@@
In this study, we define memory as an agent’s ability to change its action choice depending on the outcome of past games. By definition, this memory allows the agent to make more complex and intelligent decisions. When memory is introduced into a normal-form game, the players can achieve a wider range of strategies as the Nash equilibria (known as Folk theorem [23]). Furthermore, memory is also introduced into learning algorithms, such as replicator dynamics [24, 25, 26, 27] and Q-learning [28, 29, 30, 31, 32]. Here, since this memory causes feedback from the past, the global dynamics of such learning algorithms become more complex. Indeed, replicator dynamics diverge from the Nash equilibrium under symmetric memory lengths between players [26], while converging under asymmetric memory lengths [27]. Here, KL divergence is no longer useful to capture the global dynamics because it increases or decreases over time. The analysis of the dynamics in with-memory games is limited to the local, linearized stability analysis in the vicinity of Nash equilibrium [27]. To summarize, since memory crucially complicates learning dynamics, the global behavior of the dynamics is still unexplored.
This study provides the first theoretical analysis of the global behavior of learning in with-memory games. We assume games where their memory structure is simplest and asymmetric; One side adopts a reactive strategy that can memorize the other’s previous action [33, 34, 35, 24, 36, 25], while the other has no memory. In order to characterize the global behavior of such with-memory games, we extend KL divergence and prove that such extended divergence increases or decreases with time depending on the reactive strategy (see Fig. 1A). Furthermore, we propose a family of Lyapunov functions that characterize the dynamics of reactive strategy (see Fig. 1B). These Lyapunov functions show that the with-memory side monotonically learns to exploit the no-memory side. As an application of these functions, we prove that the convergence from arbitrary initial strategies, i.e., global convergence, occurs in matching pennies. We also experimentally confirm that such global convergence is observed in general zero-sum games.
@@
Figure 1: A. Illustration of the global behavior of the conditional divergence, $D(\bm{X},\bm{y})$. Three trajectories (red, black, and blue) are plotted with the Nash equilibrium (the black star marker). The horizontal and vertical axes show X’s strategy ($x_{1}^{{\rm st}}$) and Y’s strategy ($y_{1}$) in matching pennies (formulated in Fig. 2). This divergence decreases (red: $\dot{D}<0$), cycles (black: $\dot{D}=0$), or increases (blue: $\dot{D}>0$) with time. These three lines are plotted for the different initial strategies, i.e., $\bm{X}$ and $\bm{y}$. B. Illustration of the global behavior of the family of Lyapunov functions, $H(\bm{X};\bm{\delta})$. The colored line shows a trajectory (from purple to red) of Lyapunov functions $H_{1}$, $H_{2}$, and $H_{3}$, each of which is $H(\bm{X};\bm{\delta})$ for some specific $\bm{\delta}$. The gray broken lines are the projections of the black solid line to $H_{1}$-$H_{2}$, $H_{2}$-$H_{3}$, and $H_{3}$-$H_{1}$ planes. All of $H_{1}$, $H_{2}$, and $H_{3}$ monotonically increase with time.
2 Preliminary
2.1 Settings
@@
Figure 2: Illustration of games between reactive and zero-memory strategies. The area surrounded by the magenta dotted line shows the normal-form game. In each round, X chooses action $i=1$ or $2$ in the row, following its strategy, i.e., the probability distribution of $\bm{x}=(x_{1},x_{2})$. On the other hand, Y chooses action $j=1$ or $2$ in the column, following its strategy, i.e., the probability distribution of $\bm{y}=(y_{1},y_{2})$. Depending on their choices $i$ and $j$, X receives a payoff $u_{ij}$, given by a matrix form of $\bm{U}=(u_{ij})_{i,j}=((u_{11},u_{12}),(u_{21},u_{22}))$. Furthermore, in zero-sum games, Y receives $-u_{ij}$. Especially in the matching pennies, their actions of $1$ ($2$) correspond to the choice of “head” (“tail”) of a coin. When their choices match $i=j$, X wins, i.e., $u_{11}=u_{22}=1$ (the orange blocks). Else when their choices mismatch $i\neq j$, Y wins, i.e., $u_{12}=u_{21}=-1$ (the blue blocks). The area outside of the magenta dotted line shows the difference due to an effect of memory. The gray box shows that X memorizes Y’s previous action, represented as $j=1$ or $2$. Thus, X uses a reactive strategy and can choose its action with the conditional probability of $x_{1|j}$ and $x_{2|j}$ for Y’s previous action.
We assume that X can use reactive strategies, i.e., can change its action choice depending on the other’s previous action. This reactive strategy is denoted as $\bm{X}:=(x_{i|j})_{1\leq i\leq m_{\sf X},1\leq j\leq m_{\sf Y}}\in\prod_{1\leq j\leq m_{\sf Y}}\Delta^{m_{\sf X}-1}$, a matrix composed of $m_{\sf Y}$ vectors each of which are an element of a $m_{\sf X}-1$-dimensional simplex. Here, $x_{i|j}$ means the probability that X chooses $a_{i}$ in the condition when Y’s previous action is $b_{j}$. Thus, $\Sigma_{i}x_{i|j}=1$ should be satisfied for all $j$. On the other hand, Y only can use classical mixed strategies and choose its own action without reference to the previous actions. This mixed strategy is denoted as $\bm{y}=(y_{j})_{1\leq j\leq m_{\sf Y}}\in\Delta^{m_{\sf Y}-1}$, a vector which is an element of a $(m_{\sf Y}-1)$-dimensional simplex. Thus, $\Sigma_{j}y_{j}=1$ should be satisfied.
2.2 Stationary state and expected payoff
@@ 2.3 Nash equilibrium
$\begin{cases}\partial u^{{\rm st}}/\partial x_{i}=\Sigma_{j}u_{ij}y_{j}^{*}=C&(x_{i}^{*}>0)\\ \partial u^{{\rm st}}/\partial x_{i}=\Sigma_{j}u_{ij}y_{j}^{*}<C&(x_{i}^{*}=0)\\ \end{cases},$ (2)
$\begin{cases}\partial u^{{\rm st}}/\partial y_{j}=\Sigma_{j}u_{ij}x_{i}^{*}=C&(y_{i}^{*}>0)\\ \partial u^{{\rm st}}/\partial y_{i}=\Sigma_{j}u_{ij}x_{i}^{*}>C&(y_{i}^{*}=0)\\ \end{cases}.$ (3)
@@
2.4 Learning algorithm: replicator dynamics
@@
$\dot{x}_{i|j}$ $=x_{i|j}\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}-\Sigma_{i}x_{i|j}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}\right),$ (5)
@@
Here, following the theorems in [26], X’s and Y’s replicator dynamics include the gradient for the expected payoff $u^{{\rm st}}$. Thus, the update of X’s strategy increases its payoff $u^{{\rm st}}$, while that of Y’s strategy decreases the other’s payoff $u^{{\rm st}}$.
3 Novel quantities characterizing learning dynamics
@@
3.1 Polynomial expressions of learning dynamics
@@
Here, we remark that Eqs. (5) and (6) are nonlinear functions of $\bm{X}$ and $\bm{y}$, which is a feature of learning in with-memory games. Notably, however, these equations are polynomial expressions with $\bm{X}$ and $\bm{y}$. Such polynomial expressions cannot be seen in the games of other memory lengths [26, 27] but are special in the games between reactive and no memory strategies.
3.2 Positive definiteness for zero-sum vectors
@@
3.3 Extended Kullback-Leibler divergence
@@
If $\bm{X}^{\rm T}\bm{U}$ is positive definite for zero-sum vector, $D^{\dagger}(\bm{X};{\rm d}\bm{y}):=\dot{D}(\bm{X},\bm{y})<0$ for all ${\rm d}\bm{y}\neq\bm{0}$.
@@
where we defined ${\rm d}\bm{y}:=\bm{y}-\bm{y}^{*}$, which means the difference from Y’s equilibrium strategy and is a zero-sum vector. Thus, when $\bm{X}^{\rm T}\bm{U}$ is positive definite for zero-sum vectors, ${\rm d}\bm{y}^{\rm T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}$ is always positive, leading to $D^{\dagger}(\bm{X};{\rm d}\bm{y})<0$ for all ${\rm d}\bm{y}\neq\bm{0}$. ∎
If $\bm{X}^{\rm T}\bm{U}$ are always positive definite for zero-sum vectors, Thm. 1 gives the global behavior of learning dynamics as below.
Corollary 1 (Convergence of Y’s strategy).
If $\bm{X}^{\rm T}\bm{U}$ continues to be positive definite for zero-sum vectors, $\bm{y}=\bm{y}^{*}$ holds after sufficiently long time passes.
Proof.
This corollary is straightforwardly proved by Thm. 1. If $\bm{X}^{\rm T}\bm{U}$ is positive definite and if $\bm{y}\neq\bm{y}^{*}$, $D^{\dagger}(\bm{X};{\rm d}\bm{y})<0$ holds. Here, $D(\bm{X},\bm{y})$ has its lower bound ($\min_{\bm{X},\bm{y}}D(\bm{X},\bm{y})=0$). Thus, $D^{\dagger}(\bm{X};{\rm d}\bm{y})=0\Leftrightarrow\bm{y}=\bm{y}^{*}$ should be satisfied eventually. ∎
3.4 Family of Lyapunov functions
@@
For all $\bm{\delta}$ such that $\Sigma_{i}\delta_{i}=0$, $H(\bm{X};\bm{\delta})$ monotonically increases with time.
@@
$\dot{H}(\bm{X};\bm{\delta})=(\bm{\delta}^{\rm T}\bm{U}{\rm d}\bm{y})^{2}\ (=:H^{\dagger}(\bm{y};\bm{\delta})).$ (16)

### v2

@@ Abstract.
This study examines the global behavior of dynamics in learning in games between two players, X and Y. We consider the simplest situation for memory asymmetry between two players: X memorizes the other Y’s previous action and uses reactive strategies, while Y has no memory. Although this memory complicates their learning dynamics, we characterize the global behavior of such complex dynamics by discovering and analyzing two novel quantities. One is an extended Kullback-Leibler divergence from the Nash equilibrium, a well-known conserved quantity from previous studies. The other is a family of Lyapunov functions of X’s reactive strategy. One of the global behaviors we capture is that if X exploits Y, then their strategies converge to the Nash equilibrium. Another is that if Y’s strategy is out of equilibrium, then X becomes more exploitative with time. Consequently, we suggest global convergence to the Nash equilibrium from both aspects of theory and experiment. This study provides a novel characterization of the global behavior in learning in games through a couple of indicators.
Key words and phrases:
Multi-Agent Learning, Zero-Sum Game, Dynamical Systems, Lyapunov Function
1. Introduction
@@
Learning in games targets how multiple agents learn their optimal strategies in the repetition of games Fudenberg and Levine (1998). The set of such players’ best strategies is defined as Nash equilibrium Nash Jr (1950), where every player has no motivation to change his/her strategy. However, this equilibrium is hard to compute in general because one’s best strategy depends on the others’ strategies. Indeed, the behavior of multi-agent learning is complicated in zero-sum games, where players conflict in their payoffs. Even when players try to learn their optimal strategies there, their strategies often cycle around the Nash equilibrium and fail to converge to the equilibrium.
@@
In order to understand such strange behaviors, which are unique in multi-agent learning, the dynamics of how multiple agents learn their strategies, say, learning dynamics, are frequently studied Tuyls et al. (2006); Busoniu et al. (2008); Tuyls and Weiss (2012); Bloembergen et al. (2015). The representative dynamics of interest are the replicator dynamics, which is based on the evolutionary dynamics in biology Taylor and Jonker (1978); Friedman (1991); Börgers and Sarin (1997); Hofbauer et al. (1998); Nowak and Sigmund (2004). These dynamics are also known as the multiplicative weight updates (MWU) in its discrete-time version Arora et al. (2012); Bailey and Piliouras (2018). Furthermore, their connection to other representative learning dynamics, such as gradient ascent Singh et al. (2000); Bowling and Veloso (2002); Zinkevich (2003); Bowling (2004) and Q-learning Watkins and Dayan (1992); Hussain et al. (2023b, a), should be noted. Such replicator dynamics are known to be characterized by Kullbuck-Leibler (KL) divergence, which is the distance from the Nash equilibrium to the players’ present strategies. This KL divergence is conserved during the learning dynamics, and the distance from the Nash equilibrium is invariant Piliouras and Shamma (2014); Piliouras et al. (2014). Follow the Regularized Leader (FTRL) is a class of learning algorithms including the replicator dynamics and also has its conserved quantity, which is the summation of divergences for all the players Mertikopoulos and Sandholm (2016); Mertikopoulos et al. (2018). To summarize, such complex learning dynamics have been discussed based on their conserved quantity.
@@
In this study, we define memory as an agent’s ability to change its action choice depending on the outcome of past games. By definition, this memory allows the agent to make more complex and intelligent decisions. When memory is introduced into a normal-form game, the players can achieve a wider range of strategies as the Nash equilibria (known as Folk theorem Fudenberg and Maskin (2009)). Furthermore, memory is also introduced into learning algorithms, such as replicator dynamics Fujimoto and Kaneko (2019, 2021); Fujimoto et al. (2023, 2024) and Q-learning Masuda and Ohtsuki (2009); Barfuss (2020); Usui and Ueda (2021); Meylahn et al. (2022); Ueda (2023). Here, since this memory causes feedback from the past, the global dynamics of such learning algorithms become more complex. Indeed, replicator dynamics diverge from the Nash equilibrium under symmetric memory lengths between players Fujimoto et al. (2023), while converging under asymmetric memory lengths Fujimoto et al. (2024). Here, KL divergence is no longer useful to capture the global dynamics because it increases or decreases over time. The analysis of the dynamics in with-memory games is limited to the local, linearized stability analysis in the vicinity of Nash equilibrium Fujimoto et al. (2024). To summarize, since memory crucially complicates learning dynamics, the global behavior of the dynamics is still unexplored.
@@
Figure 1. (A) Illustration of the global behavior of the conditional divergence, $D(\bm{X},\bm{y})$. Three trajectories (red, black, and blue) are plotted with the Nash equilibrium (the black star marker). The horizontal and vertical axes show X’s strategy ($x_{1}^{{\rm st}}$) and Y’s strategy ($y_{1}$) in the matching pennies game (formulated in Fig. 2). This divergence decreases (red: $\dot{D}<0$), cycles (black: $\dot{D}=0$), or increases (blue: $\dot{D}>0$) with time. These three lines are plotted for the different initial strategies, i.e., $\bm{X}$ and $\bm{y}$. (B) Illustration of the global behavior of the family of Lyapunov functions, $H(\bm{X};\bm{\delta})$. The colored line shows a trajectory (from purple to red) of Lyapunov functions $H_{1}$, $H_{2}$, and $H_{3}$, each of which is $H(\bm{X};\bm{\delta})$ for some specific $\bm{\delta}$. The gray broken lines are the projections of the black solid line to $H_{1}$-$H_{2}$, $H_{2}$-$H_{3}$, and $H_{3}$-$H_{1}$ planes. All of $H_{1}$, $H_{2}$, and $H_{3}$ monotonically increase with time.
This study provides the first theoretical analysis of the global behavior of learning in with-memory games. We assume games where their memory structure is simplest and asymmetric; One side adopts a reactive strategy that can memorize the other’s previous action Nowak (1990); Ohtsuki (2004); Baek et al. (2016); Fujimoto and Kaneko (2019); Schmid et al. (2022); Fujimoto and Kaneko (2021), while the other has no memory. In order to characterize the global behavior of such with-memory games, we extend KL divergence and prove that such extended divergence increases or decreases with time depending on whether the reactive strategy is exploitative or not (see Fig. 1A). We further propose a family of Lyapunov functions that characterize the dynamics of the reactive strategy (see Fig. 1B). These Lyapunov functions show that the with-memory side monotonically learns to exploit the no-memory side. As an application of these functions, we suggest the convergence from arbitrary initial strategies to the equilibrium, i.e., global convergence. We prove global convergence in the matching pennies game. We also experimentally confirm that such global convergence is observed in other games equipped with various types of equilibrium.
2. Preliminary
2.1. Settings
@@
Figure 2. Illustration of games between reactive and zero-memory strategies. The area surrounded by the magenta dotted line shows the normal-form game. In each round, X chooses action $i=1$ or $2$ in the row, following its strategy, i.e., the probability distribution of $\bm{x}=(x_{1},x_{2})$. On the other hand, Y chooses action $j=1$ or $2$ in the column, following its strategy, i.e., the probability distribution of $\bm{y}=(y_{1},y_{2})$. Depending on their choices $i$ and $j$, X receives a payoff $u_{ij}$, given by a matrix form of $\bm{U}=(u_{ij})_{i,j}=((u_{11},u_{12}),(u_{21},u_{22}))$. Furthermore, in zero-sum games, Y receives $-u_{ij}$. Especially in the matching pennies game, their actions of $1$ ($2$) correspond to the choice of “head” (“tail”) of a coin. When their choices match $i=j$, X wins, i.e., $u_{11}=u_{22}=1$ (the orange blocks). Else when their choices mismatch $i\neq j$, Y wins, i.e., $u_{12}=u_{21}=-1$ (the blue blocks). The area outside of the magenta dotted line shows the difference due to an effect of memory. The gray box shows that X memorizes Y’s previous action, represented as $j=1$ or $2$. Thus, X uses a reactive strategy and can choose its action with the conditional probability of $x_{1|j}$ and $x_{2|j}$ for Y’s previous action.
We assume that X use reactive strategies, i.e., can change its action choice depending on the other’s previous action. This reactive strategy is denoted as $\bm{X}:=(x_{i|j})_{1\leq i\leq m_{\sf X},1\leq j\leq m_{\sf Y}}\in\prod_{1\leq j\leq m_{\sf Y}}\Delta^{m_{\sf X}-1}$, a matrix composed of $m_{\sf Y}$ vectors each of which are an element of a $m_{\sf X}-1$-dimensional simplex. Here, $x_{i|j}$ means the probability that X chooses $a_{i}$ in the condition when Y’s previous action is $b_{j}$. Thus, $\Sigma_{i}x_{i|j}=1$ should be satisfied for all $j$. On the other hand, Y only can use classical mixed strategies and choose its own action without reference to the previous actions. This mixed strategy is denoted as $\bm{y}=(y_{j})_{1\leq j\leq m_{\sf Y}}\in\Delta^{m_{\sf Y}-1}$, a vector which is an element of a $(m_{\sf Y}-1)$-dimensional simplex. Thus, $\Sigma_{j}y_{j}=1$ should be satisfied.
2.2. Stationary State and Expected Payoff
@@ 2.3. Nash Equilibrium
$\begin{cases}\partial u^{{\rm st}}/\partial x_{i}=\Sigma_{j}u_{ij}y_{j}^{*}=C&(x_{i}^{*}>0)\\ \partial u^{{\rm st}}/\partial x_{i}=\Sigma_{j}u_{ij}y_{j}^{*}\leq C&(x_{i}^{*}=0)\\ \end{cases},$ (2)
$\begin{cases}\partial u^{{\rm st}}/\partial y_{j}=\Sigma_{i}u_{ij}x_{i}^{*}=C&(y_{j}^{*}>0)\\ \partial u^{{\rm st}}/\partial y_{j}=\Sigma_{i}u_{ij}x_{i}^{*}\geq C&(y_{j}^{*}=0)\\ \end{cases}.$ (3)
@@
2.4. Learning Algorithm: Replicator Dynamics
@@
$\dot{x}_{i|j}$ $=+x_{i|j}\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}-\Sigma_{i}x_{i|j}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}\right),$ (5)
@@
Here, following the theorems in Fujimoto et al. (2023), X’s and Y’s replicator dynamics include the gradient for the expected payoff $u^{{\rm st}}$. Thus, the update of X’s strategy increases its payoff $u^{{\rm st}}$, while that of Y’s strategy decreases the other’s payoff $u^{{\rm st}}$. We discuss learning based on the replicator dynamics throughout this study, but we can extend all the following results to another typical learning algorithm, the gradient descent-ascent (see Appendix D for detailed discussion).
3. Theory on Learning Dynamics
@@
3.1. Polynomial Expressions of Learning
@@
Here, we remark that Eqs. (5) and (6) are nonlinear functions of $\bm{X}$ and $\bm{y}$, which is a feature of learning in with-memory games. Notably, however, these equations are polynomial expressions with $\bm{X}$ and $\bm{y}$. Such polynomial expressions cannot be seen in the games of other memory lengths Fujimoto et al. (2023, 2024) but are special in the games between reactive and no memory strategies.
3.2. Positive Definiteness for Zero-Sum Vectors
@@
3.3. Extended Kullback-Leibler Divergence
@@
If $\bm{X}^{\rm T}\bm{U}$ is positive definite for zero-sum vector, $D^{\dagger}(\bm{X};{\rm d}\bm{y}):=\dot{D}(\bm{X},\bm{y})<0$ for all ${\rm d}\bm{y}:=\bm{y}-\bm{y}^{*}\neq\bm{0}$.
@@
which means the difference from Y’s equilibrium strategy and is a zero-sum vector. Thus, when $\bm{X}^{\rm T}\bm{U}$ is positive definite for zero-sum vectors, ${\rm d}\bm{y}^{\rm T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}$ is always positive, leading to $D^{\dagger}(\bm{X};{\rm d}\bm{y})<0$ for all ${\rm d}\bm{y}\neq\bm{0}$. ∎
3.4. Family of Lyapunov Functions
@@
For all $\bm{\delta}$ such that $\Sigma_{i}\delta_{i}=0$, $H^{\dagger}(\bm{y};\bm{\delta}):=\dot{H}(\bm{X};\bm{\delta})\geq 0$. The equality holds if and only if $d\bm{y}(=\bm{y}-\bm{y}^{*})=\bm{0}$.
@@
$\dot{H}(\bm{X};\bm{\delta})=|\bm{\delta}^{\rm T}\bm{U}{\rm d}\bm{y}|^{2}\ (=:H^{\dagger}(\bm{y};\bm{\delta})).$ (16)